With what minimum speed must you toss a 160g ball straight up to hit the 13m -high roof of the gymnasium if you release the ball 1.4m above the ground? Solve this problem using energy. With what speed does the ball hit the ground?
Answer #1


apply KE = chnage of PE

0.5 m V^2 = mg (h2-h1)

V^2 = 2* 9.81*(13-1.4)

V = 15.08 m/s
Answer #2

1. energy conservation

1/2*m*v^2 + mgh1 = mgh2

v = sqrt(2*g*(h2-h1)) = 15.08m/s

when it hits ground

1/2*m*15.06^2 + m*9.8*1.4 = 1/2*m*v^2

v = 15.94m/s